QUESTION 1

  1. Anglo Foods Inc., has the following forecasted sales for 2018:

    January

    February

    March

    April

    May

    June

    July

    $200,000

    $350,000

    $420,000

    $350,000

    $270,000

    $200,000

    $300,000

    Actual sales in November and December 2017 were $375,000 and 266,667, respectively. 60% of sales are on credit (i.e. 40% of sales will be received in cash in the same month). The firm collects 60% of these credit sales (i.e., 60%*60% = 36% of sales) during the first month after the sale and the remainder (i.e. 40% of credit sales and thus 40%*60% = 24% of sales) during the following second month.

    For example, for the January sales of $200,000, 40% (i.e. $80,000) will be received in cash in January, 36% (i.e. $72,000) will be received in cash in February, and the rest ($48,000) will be received in cash in March.

           Here are the information about the cash expenditures.

    1. Purchases constitute 60% of the next month’s sales. The company pays 50% during the first month after the purchase was made and the remainder in the next month.

    For example, January purchases will be 60% of February sales, i.e., $350,000*60%=$210,000. The company pays 50% of it ($105,000) in cash in February, and the rest in March.

    1. Wages, taxes and other expenses are expected to be 30% of forecasted sales.
    2. A major capital expenditure of $70,000 is expected in March.
    3. Interest payments are expected to be $30,000 every month.

    The company needs a minimum cash balance of $25,000. The beginning cash balance is $30,000.

    Using the Bithlo Barbecues’ simple cash budget example from Chapter 4, prepare their cash budget for January to June 2018.

    Answer Q (1-5) based on the cash budget you have prepared.

    1. How much is the company’s cash collection in March from the March sales of $420,000?
    a

    $168,000

    b

    $126,000

    c

    $342,000

    d

    $400,000
Answer #1

The cash collection of March sales will be as follows :

Amount
cash sales 420000*40%= 168000
credit sales
April 420000*36%= 151200
May 420000*24%= 100800
Total 420000

company’s cash collection in March from the March sales of $420,000 = $ 168000

correct option is "a"- 168000